# NOBLE - Neural Operator with Biologically-informed Latent Embeddings to Capture Experimental Variability in Biological Neuron Models

### v2

@@ Abstract
Characterizing the diverse computational properties of human neurons via multimodal electrophysiological, transcriptomic, and morphological data provides the foundation for constructing and validating bio-realistic neuron models that can advance our understanding of fundamental mechanisms underlying brain function. However, current modeling approaches remain constrained by the limited availability and intrinsic variability of experimental neuronal data. To capture variability, ensembles of deterministic models are often used, but are difficult to scale as model generation requires repeating computationally expensive optimization for each neuron. While deep learning is becoming increasingly relevant in this space, it fails to capture the full biophysical complexity of neurons, their nonlinear voltage dynamics, and variability. To address these shortcomings, we introduce NOBLE, a neural operator framework that learns a mapping from a continuous frequency-modulated embedding of interpretable neuron features to the somatic voltage response induced by current injection. Trained on data generated from biophysically realistic neuron models, NOBLE predicts distributions of neural dynamics accounting for the intrinsic experimental variability. Unlike conventional bio-realistic neuron models, interpolating within the embedding space offers models whose dynamics are consistent with experimentally observed responses. NOBLE is the first scaled-up deep learning framework validated on real experimental data, enabling efficient generation of synthetic neurons that exhibit trial-to-trial variability and achieve a 4200×4200$\times$4200 × speedup over numerical solvers. To this end, NOBLE captures fundamental neural properties, opening the door to a better understanding of cellular composition and computations, neuromorphic architectures, large-scale brain circuits, and general neuroAI applications.
1Department of Mathematics, ETH Zürich, Zürich, Switzerland
2Department of Computing and Mathematical Sciences, California Institute of Technology, Pasadena, CA, USA
3Department of Electrical and Computer Engineering, University of Alberta, Alberta, Canada
4Neuroscience and Mental Health Institute, University of Alberta, Alberta, Canada
5Department of Neurosurgery, Cedars-Sinai Medical Center, Los Angeles, CA, USA
6Department of Neurology, Cedars-Sinai Medical Center, Los Angeles, CA, USA
7Center for Biomedical Science, Cedars-Sinai Medical Center, Los Angeles, CA, USA
8Archimedes AI, Athena Research Center, Marousi, Greece
@@ 1 Introduction
Hundreds of distinct neuronal cell types co-exist and compute within neural circuits, yet how they shape cognitive functions remains essentially unanswered [1, 2, 3, 4, 5]. This is particularly true in the human brain, where access and monitoring capabilities are severely limited compared to animal models. Over the past decade, multimodal cellular datasets that integrate electrophysiology, morphology, and transcriptomics have emerged for human cell types [6, 7, 8, 9, 10]. While integrating across the different modalities remains a challenge, clear differences in gene expression, morphology, and electrophysiology are evident across cell types. However, how such diversity scales up to circuit-level function is still unknown. Understanding how these differences affect overall brain function is crucial in many contexts, for instance, to uncover genes or phenotypes related to neurological diseases.
@@
Cellular models representing multiple data modalities are invaluable as they offer a degree of control and perturbations that are experimentally infeasible (e.g., [11, 12, 13]). Recently, evolutionary multi-objective optimization algorithms [14] have been used to generate and validate realistic models of neurons in the form of 3D multi-compartment partial differential equation (PDE) models that mirror both their shape and ion channel expression, shaping their electrical properties [12, 15, 13, 16]. The downside is that such models are deterministic and fail to capture the variability observed experimentally, where identical input to the same neuron often results in different electrophysiological responses. One approach is to generate families of models, sometimes referred to as "hall-of-fame" (HoF) models [12, 13, 17] to represent a single cell.
While each HoF model is distinct and reproduces the electrophysiological features of parts of an experiment, the ensemble of deterministic HoF models is used as a collective representation that captures both the main features as well as their variability in an experiment [12, 13]. Typically, neurons exhibit highly nonlinear behavior, necessitating equally complex models with a large number of parameters, which makes multi-objective evolutionary optimization computationally demanding, requiring about 600k CPU core hours per single-neuron model [12, 13]. Yet, even tiny perturbations of the model parameters lead to large deviations from experimental data [18]. In summary, capturing the nature and variability of neurons is a challenge with existing computational techniques.
@@
The challenges of scalability and computational cost associated with traditional modeling approaches, such as numerical integrators and evolutionary optimization algorithms, have led the scientific community to explore the use of machine learning to accelerate simulations by learning underlying relationships between variables directly from data. While neural networks have been used successfully for many applications, they learn mappings between finite-dimensional vectors, which can limit their ability to model physical phenomena that are better described using functions in infinite-dimensional spaces and functional relationships between them [19]. As a result, neural networks can overfit to the training discretization and suffer from limited out-of-distribution capabilities. Neural operators [19, 20] are a principled way to generalize neural networks to learn operators mapping functions to functions, with a universal operator approximation property [21]. A variety of neural operators have been proposed, such as the Fourier Neural Operator (FNO) [22], and successfully applied to a wide range of problems [23, 24, 25]. Prior knowledge of the relevant physics laws and differential equations can also be incorporated as additional loss terms during training, to supplement or replace reference data, as done with physics-informed neural operators [26, 27].
Earlier applications of machine learning to single-cell electrophysiology focused on directly learning the dynamics of canonical point-neuron models like FitzHugh-Nagumo (FHN) [28, 29] and Hodgkin-Huxley (HH) [30]. Fully connected and convolutional neural networks were trained to reproduce FHN dynamics [31], while ResNet-based multilayer perceptrons showed promise in learning HH dynamics [32, 33]. Physics-informed neural networks (PINNs) extend conventional neural networks by introducing prior knowledge about the underlying dynamical system [34, 35, 36]. This formulation was used to predict FHN dynamics [37] and to learn the HH model ionic conductances from simulated voltage recordings [38]. Further refinements with PINNs incorporated wavelet bases to capture localized multiscale dynamics and compute derivatives analytically, improving both accuracy and convergence speed when training on the FHN model [39]. While these methods demonstrate capabilities in capturing the core spiking dynamics of these simplified models, their ability to accurately represent the full spectrum of electrophysiological behavior, particularly the highly nonlinear onset of firing, remains largely untested. Moreover, as function approximators, they necessitate retraining for each new input stimulus, significantly limiting their practical utility.
To address these limitations, Centofanti et al. [40] explored using operator learning approaches for forward simulations of the HH model. Among other approaches, FNOs showed promising results by demonstrating a strong capacity for learning the governing operators of this biophysical system. However, this work still exhibits key limitations and a limited scope: (1) it relies on relatively simple simulated data from a point-neuron model, (2) it does not explicitly attempt to capture the full spectrum of electrophysiological dynamics, particularly the highly nonlinear onset of firing, and (3) its formulation on a single operator inherently lacks the capacity to represent the trial-to-trial variability observed in biological recordings. This paper builds on the promise of operator learning for neuroscience demonstrated in [40], extending its scope and addressing its limitations to enable deeper insights into brain function and its applications in neuroAI.
@@
Figure 1: The Neural Operator with Biologically-informed Latent Embeddings (NOBLE) framework. A) In NOBLE, a current injection and neuron model features are first encoded using the proposed embedding strategy, before passing through a neural operator to produce a prediction for the somatic voltage response. B) NOBLE can be queried in parallel with different model latent representations to produce ensemble predictions. C) The proposed embedding in NOBLE encodes specified neuron features and the input current as a stack of trigonometric time-series, as described in Section 3.3.
@@
Contributions.  We introduce NOBLE (Neural Operator with Biologically-informed Latent Embeddings), a neural operator framework for learning the nonlinear somatic dynamics across a population of HoF models for a single neuron (Figure 1). NOBLE is the first scaled-up deep learning framework whose performance is validated with experimental data. Rather than training a separate independent surrogate for each HoF model, NOBLE learns a single neural operator that maps from a continuous latent space of user-defined, interpretable neuron characteristics to an ensemble of somatic voltage responses induced by current injection. This latent space is constructed using an embedding strategy informed by the specified characteristics of the neuron models.
@@
As an example application, we instantiate NOBLE with an FNO, and train and evaluate it on a parvalbumin-positive (PVALB) neuron dataset generated using 50 HoF PVALB models (Figure 2). We show that a single NOBLE model accurately captures both subthreshold and spiking dynamics across all 50 HoF models (in-distribution) as well as unseen HoF (out-of-distribution) while achieving a significant speedup of $4200$× over the numerical solver used to generate the dataset (Figure 4). In addition, the NOBLE predictions across 15 electrophysiological features of interest (including spike count, amplitude, and width) remain within the variability observed in experimental data (Figure 7). We further show that NOBLE can successfully generate novel bio-realistic neuron models by sampling and interpolating within the latent space of models. The dynamics of novel neuron models generated by NOBLE align both with previously unseen HoF PVALB models and experimentally observed somatic responses. In contrast, direct interpolation between the parameters of bio-realistic PDE-based neuron models fails due to the sensitivity and nonlinearity of the underlying PDEs [41, 42] (Figure 5).
@@
Owing to NOBLE’s ability to generate novel bio-realistic neuron models, ensemble predictions are no longer constrained to the original 50 HoF models used during training. NOBLE can produce somatic voltage responses for an arbitrary number of biologically plausible neurons. We demonstrate this enhanced flexibility by predicting responses to input stimuli across a larger set of models (Figure 6). The results showcase NOBLE’s ability to accurately capture a broad range of biophysically realistic neurons while enabling dense interpolation across model space. By unlocking the efficient generation of infinitely many diverse yet biophysically plausible neurons from a continuous embedding, NOBLE offers a scalable alternative to computationally intensive, scale-limited evolutionary approaches, laying the foundation for brain-scale neural circuit modeling.
@@
Finally, the biologically-informed latent representation of the neuron models together with the capability of NOBLE to generate arbitrarily many new biophysically-realistic neuron models also offers further insight into the behavior of neural dynamics. We can use NOBLE to obtain the somatic responses to current injections on a fine grid in the model latent representation space, and consequently construct heat maps and surface plots to better understand how neuron features used for the model latent representation affect any electrophysiological feature of interest.
2 Background
2.1 Bio-realistic Neuron Modeling
We create bio-realistic PDE-based neuron models using actual reconstructions of neuron morphologies from human cortical data [6, 8, 9] (Figure 2). We instantiate these models using a framework [41] built on the NEURON simulation environment [42]. We place ion channels in an "all-active" configuration [12, 13], where active ion channels are distributed along both somatic and dendritic compartments along the neuron morphology. For each experimental neuron, models are generated using a multi-objective evolutionary optimization framework [13] to find ion conductance parameters replicating a standard set of electrophysiological features from patch clamp recordings (Figure 2A). We adopt a two-stage optimization strategy, first fitting passive subthreshold responses followed by capturing the active dynamics above the spiking threshold and the full frequency-current curve of each neuron. After 250 generations of evolutionary optimization, the models that best minimize the mean z-score error between simulated and actual experimental electrophysiological features are selected as hall-of-fame (HoF) models (Figure 2B). More details about the electrophysiological features of interest are provided in Appendix E.
@@ 2.1 Bio-realistic Neuron Modeling
To illustrate the proposed approach, we consider in this paper a randomly selected PVALB human cortical neuron, for which we have created 60 HoF models. PVALB neurons are fast-spiking inhibitory interneurons regulating high frequency gamma oscillations (30-80Hz) [43] and their dysfunction has been associated with cognitive impairments such as schizophrenia and Alzheimer’s disease [44, 45].
2.2 Neural Operators
Neural operators compose linear integral operators $\mathcal{K}$ with pointwise nonlinear activation functions $\sigma$ to approximate highly nonlinear operators. More precisely, we define the neural operator
where $\mathcal{P}$, $\mathcal{Q}$ are the pointwise neural networks that encode the lower dimension function into a higher dimensional space and vice versa. The model stacks $L$ layers of $\sigma(W_{l}+\mathcal{K}_{l}+b_{l})$ where $W_{l}$ are pointwise linear operators (matrices), $\mathcal{K}_{l}$ are integral kernel operators, $b_{l}$ are bias terms, and $\sigma$ are fixed activation functions. The parameters $\theta$ consists of all the parameters in $\mathcal{P},\mathcal{Q},W_{l},\mathcal{K}_{l},b_{l}$. Kossaifi et al. [46] maintain a comprehensive open-source PyTorch library for learning neural operators, which serves as the foundation for our implementation.
A Fourier neural operator (FNO) [22] is a neural operator which uses Fourier integral operator layers $\bigl{(}\mathcal{K}(\phi)v_{t}\bigr{)}(x)=\mathcal{F}^{-1}\Bigl{(}R_{\phi}\cdot(\mathcal{F}v_{t})\Bigr{)}(x)$ where $R_{\phi}$ is the Fourier transform of a periodic function $\kappa$ parameterized by $\phi$. On a uniform mesh, the Fourier transform $\mathcal{F}$ can be implemented using the fast Fourier transform (FFT). A depiction of the FNO architecture is provided in Figure 8.
@@ 3 Method
We present the proposed NOBLE framework, along with the additional data generation and processing mechanisms we developed to enhance training efficiency and reduce computational cost. Specifically, we begin in Section 3.1 by introducing a subsampling strategy for somatic voltage responses that preserves the key neuronal features of interest. We then propose a non-uniform training distribution of current injection amplitudes in Section 3.2 to effectively capture the transition from non-spiking to spiking behavior. Next, we present the neuron model embedding strategy at the core of NOBLE in Section 3.3, and conclude by detailing the full architecture of NOBLE in Section 3.4.
@@ 3.1 Subsampling
NOBLE employs a notable property of neural operators on training on low-resolution data while reserving the capability to generate dynamics at higher resolution. In this regard, we subsample the reference HoF simulations in time. To avoid discarding high-resolution information necessary for capturing neuron features of interest, we analyze how these features are affected by different subsampling factors and strategies. In particular, we compare their effect on the discrepancy between HoF simulations and experimental data. We consider (i) low-pass filtering followed by decimation in time, (ii) low-pass filtering followed by truncation in the frequency domain, (iii) direct truncation in the frequency domain, and (iv) decimation in time without filtering. Across neuron features, we observe no consistent differences in performance between these strategies and thus opt for low-pass filtering followed by decimation in time. Our analysis (see Appendix B) reveals that a subsampling factor of 3 preserves the fidelity of extracted features, introducing no more error than the intrinsic discrepancy between HoF simulations and experimental data, and does not lead to significant aliasing. This reduces the time series length from $25,\!750$ to $8,\!583$, substantially decreasing the computational load without compromising biological realism.

### v3

# NOBLE – Neural Operator with Biologically-informed Latent Embeddings to Capture Experimental Variability in Biological Neuron Models

@@ Abstract
Characterizing the cellular properties of neurons is fundamental to understanding their function in the brain. In this quest, the generation of bio-realistic models is central towards integrating multimodal cellular data sets and establishing causal relationships. However, current modeling approaches remain constrained by the limited availability and intrinsic variability of experimental neuronal data. The deterministic formalism of bio-realistic models currently precludes accounting for the natural variability observed experimentally. While deep learning is becoming increasingly relevant in this space, it fails to capture the full biophysical complexity of neurons, their nonlinear voltage dynamics, and variability. To address these shortcomings, we introduce NOBLE, a neural operator framework that learns a mapping from a continuous frequency-modulated embedding of interpretable neuron features to the somatic voltage response induced by current injection. Trained on synthetic data generated from bio-realistic neuron models, NOBLE predicts distributions of neural dynamics accounting for the intrinsic experimental variability. Unlike conventional bio-realistic neuron models, interpolating within the embedding space offers models whose dynamics are consistent with experimentally observed responses. NOBLE enables the efficient generation of synthetic neurons that closely resemble experimental data and exhibit trial-to-trial variability, offering a 42004200× speedup over the numerical solver. NOBLE is the first scaled-up deep learning framework that validates its generalization with real experimental data. To this end, NOBLE captures fundamental neural properties in a unique and emergent manner that opens the door to a better understanding of cellular composition and computations, neuromorphic architectures, large-scale brain circuits, and general neuroAI applications.
@@ 1 Introduction
Hundreds of distinct neuronal cell types co-exist and compute within neural circuits, yet how they shape cognitive functions remains essentially unanswered [1, 2, 3, 4, 5]. This is particularly true in the human brain, where access and monitoring capabilities are severely limited compared to animal models. Over the past decade, multimodal cellular datasets that integrate electrophysiology, morphology, and transcriptomics have emerged for human cell types [6, 7, 8, 9, 10].
While integrating across the different modalities remains a challenge, clear differences in gene expression, morphology, and electrophysiology are evident across cell types. However, understanding how these differences impact brain processing is crucial, e.g. to uncover how expression of specific genes relates to neurological diseases.
@@
Cellular models representing multiple data modalities are invaluable as they offer a degree of control and perturbations that are experimentally impossible (e.g., [11, 12, 13]). Recently, evolutionary multi-objective optimization algorithms [14] have been used to generate and validate bio-realistic models of neurons in the form of 3D multi-compartment partial differential equation (PDE) models that mirror both their shape and ion channel expression, shaping their electrical properties [12, 13, 15, 16]. Yet, such models are deterministic and fail to capture the intrinsic variability observed experimentally, where identical input to the same neuron often results in different electrophysiological responses. One approach is to generate families of models, referred to as "hall-of-fame" (HoF) models [12, 13, 17] to represent a single cell. While each HoF model is distinct and reproduces the electrophysiological features of parts of an experiment, the ensemble of deterministic HoF models is used as a collective representation that captures both the main features as well as their variability in an experiment [12, 13]. Typically, neurons exhibit highly nonlinear behavior, necessitating equally complex models rendering the optimization computationally demanding (i.e. requiring about 600k CPU core hours per single-neuron model [12, 13]). Yet, even tiny perturbations of the model parameters lead to large deviations from experimental data [18]. Other approaches have explored capturing variability through introducing stochasticity in neuron models [19, 20, 21]. However, the synthetic injection of white noise is non-mechanistic and introduces perturbations that can lead to unrealistic predictions [22, 23, 24]. In summary, capturing the nature and variability of neurons is a challenge with existing computational techniques.
@@
The challenges of scalability and the computational cost associated with traditional numerical modeling approaches, such as numerical integrators and evolutionary optimization algorithms, have led the scientific community to explore the use of machine learning to accelerate simulations by learning underlying relationships between variables directly from experimental and synthetic data. While neural networks have been used successfully for many applications, they learn mappings between finite-dimensional vectors, which can limit their ability to model physical phenomena that are better described using functions in infinite-dimensional spaces and functional relationships between them [25]. As a result, neural networks can overfit to the training discretization and suffer from limited out-of-distribution capabilities. Neural operators [25, 26, 27] are a principled way to generalize neural networks to learn operators mapping functions to functions, with a universal operator approximation property [28]. A variety of neural operators have been proposed, such as the Fourier Neural Operator (FNO) [29].
Machine learning approaches have been applied to model single-cell electrophysiology, primarily for point-neuron systems such as FitzHugh–Nagumo [30, 31] and Hodgkin–Huxley [32]. Conventional neural networks [33, 34, 35] and physics‑informed neural networks [36, 37, 38, 39, 40, 41] successfully reproduced their dynamics but remain highly specific to deterministic formulations and require retraining for each new stimulus. More recently, neural operators demonstrated strong potential for learning the governing dynamics of Hodgkin–Huxley systems [42], though the study was limited to simplified data, without capturing biological variability. Related works like NeuPRINT [43] captured biological neuronal variability, but on slower in vivo 2-photon calcium imaging data and models population-level fluorescence dynamics, rather than fast intracellular voltage dynamics of individual neurons.
We build on these advances and address current limitations to enable deeper insights into brain function and neuroAI.
@@
Figure 1: The Neural Operator with Biologically-informed Latent Embeddings (NOBLE) framework. A) In NOBLE, a current injection and neuron model features are first encoded using the proposed embedding strategy, before passing through a neural operator to produce a prediction for the somatic voltage response. B) NOBLE can be queried in parallel with different model latent representations to produce ensemble predictions. C) The proposed embedding in NOBLE encodes specified neuron features and the input current as a stack of trigonometric time-series, as described in Section˜3.4.
@@
Contributions. We introduce NOBLE (Neural Operator with Biologically-informed Latent Embeddings), a neural operator framework for learning the nonlinear somatic dynamics across a population of HoF models for a single neuron (Figure˜1). NOBLE is the first scaled-up deep learning framework whose performance is validated with experimental human cortex data. Rather than training a separate independent surrogate for each HoF model, NOBLE learns a single neural operator that maps from a continuous latent space of user-defined, interpretable neuron characteristics to an ensemble of somatic voltage responses induced by current injection. This latent space is constructed using an embedding strategy informed by the specified characteristics of the neuron models.
@@
As an example application, we train and evaluate NOBLE on a parvalbumin-positive (PVALB) neuron dataset generated using 50 HoF PVALB models (Figure˜2). We show that a single NOBLE model accurately captures both subthreshold and spiking dynamics across all 50 HoF models (in-distribution) as well as 10 unseen HoF models (out-of-distribution) while achieving a significant speedup of 42004200× over the numerical solver used to generate the dataset (Figure˜3). In addition, the NOBLE predictions across 16 electrophysiological features of interest (including spike count, amplitude, and width) remain within the variability observed in experimental data (Figure˜5B). Additional ablation studies confirm that biologically informed embeddings are critical for capturing both firing and non-firing dynamics, and demonstrate that we can enhance performance on targeted features without compromising overall dynamics by introducing a feature-specific fine-tuning approach.
@@
We further show that NOBLE can successfully generate novel bio-realistic neuron models by sampling and interpolating within the latent space of models. The dynamics of novel neuron models generated by NOBLE align both with previously unseen HoF PVALB models and experimentally observed somatic responses. In contrast, direct interpolation between the parameters of bio-realistic PDE-based neuron models fails due to the sensitivity and nonlinearity of the underlying PDEs [44, 45] (Figure˜4). We also successfully instantiate and train an additional NOBLE on a vasoactive intestinal peptide (VIP) interneuron to demonstrate the generalizability of our embedding framework. Owing to NOBLE’s ability to generate novel bio-realistic neuron models, ensemble predictions are no longer constrained to the original 50 HoF models used for training. We demonstrate that NOBLE can produce somatic voltage responses for an arbitrary number of biologically plausible neurons by predicting responses to input stimuli across a larger set of models (Figure˜5). The results showcase NOBLE’s ability to accurately capture a broad range of neuron dynamics while enabling dense interpolation across model space. By unlocking the efficient, unlimited generation of diverse yet realistic neurons from a continuous embedding, NOBLE offers a scalable alternative to computationally intensive, scale-limited evolutionary approaches, laying the foundation for brain-scale neural circuit modeling.
@@
Finally, the biologically-informed latent representation of the neuron models together with the capability of NOBLE to generate arbitrarily many new bio-realistic neuron models also offers further insight into the behavior of neural dynamics. We can use NOBLE to obtain the somatic responses to current injections on a fine grid in the model latent representation space, and consequently construct heat maps and surface plots to better understand how neuron features used for the model latent representation affect any electrophysiological feature of interest.
2 Background on Bio-Realistic Neuron Modeling
@@
We create bio-realistic PDE-based neuron models (based on the cable equation [46]) using actual reconstructions of neuron morphologies from human cortical data [6, 8, 9] (Figure˜2). We instantiate these models using a framework [44] built on the NEURON simulation environment [45], which uses a spatial discretization to simulate the models as a system of coupled ordinary differential equations. We place ion channels in an “all-active" configuration [12, 13], where active ion channels are distributed along both somatic and dendritic compartments along the neuron morphology. For each experimental neuron, models are generated using a multi-objective evolutionary optimization framework [13] to find ion conductance parameters replicating a standard set of electrophysiological features from patch clamp recordings (Figure˜2A). We adopt a two-stage optimization strategy, first fitting passive subthreshold responses, followed by capturing the active dynamics above the spiking threshold and the full frequency-current curve of each neuron. After 250 generations of evolutionary optimization, the models that best minimize the mean z-score error between simulated and actual experimental electrophysiological features are selected as HoF models (Figure˜2B). More details about the electrophysiological features of interest are provided in Appendix˜B.
To illustrate the proposed approach, we consider a randomly selected PVALB human cortical neuron, for which we created 60 HoF models. PVALB neurons are fast-spiking inhibitory interneurons regulating high-frequency gamma oscillations (30-80Hz) [47] and their dysfunction has been associated with cognitive impairments such as schizophrenia and Alzheimer’s disease [48, 49]. We also consider the class VIP of inhibitory interneurons, known for its disinhibitory role in cortical circuits [50, 51, 52].
@@ 3.1 Subsampling
NOBLE utilizes the notable property of neural operators of training on low-resolution data while reserving the capability to generate dynamics at higher resolution. In this regard, we subsample the reference HoF simulations in time. To avoid discarding high-resolution information necessary for capturing neuron features of interest, we analyze how these features are affected by different subsampling factors and strategies, in particular via the discrepancy between HoF simulations and experimental data. We consider (i) low-pass filtering followed by decimation in time, (ii) low-pass filtering followed by truncation in the frequency domain, (iii) truncation in the frequency domain, and (iv) decimation in time without filtering. Across neuron features, we observe no consistent differences in performance between these strategies and thus opt for low-pass filtering followed by decimation in time. Our analysis (see Section˜C.1) reveals that 3× subsampling preserves the fidelity of extracted features, within the bounds of the intrinsic discrepancy between HoF simulations and experimental data, and without inducing notable aliasing. For the HoF simulations, we consider time series of $515\mathrm{ms}$ with a timestep of $0.02\mathrm{ms}$. For such signals, the subsampling reduces the sequence length from 25,75025,\!750 to 8,5838,\!583, substantially decreasing the computational load without compromising biological realism.